Bringing mathematical and physical manuscripts to the net

Dirk Wintergrün

version 0.1 of August 28, 2003

Abstract

Proposal for a joint task force/project of the Dibner Institute and the Max Planck Institute for the History of Science

Contents

1 Starting point

Both institutes aim to bring a.s.a.p. mathematical papers into the net, i.e. Albert Einstein’s Zurich Notebook by the MPIWG and Manuscripts of Newton and contemporaries by the Dibner Institute. Furthermore these projects shall be realized without substantial external funding. Therefore a strategy for using already existing technology has to be found which also includes an optimal devision of labor.

The problem of presenting images and (pure) text is to great extent solved. But an open question is how to adequately represent formulae.

2 overall goal

The overall goal of the project/task force to develop a.s.a p. an environment for presenting mathematical and physical manuscripts in the net. In particular a solution for the representation of formulae in the historic context should be found. The developments shall base on existing technologies already developed at the participating institutes and otherwise available open source technologies.

3 goals

In detail in a first step (pragmatic) answers to the following questions should be found:

• Which technologies exists for presenting formulae in the net?

• Which standards for encoding formulae as XML exists and can they be used for representing formulae in historic papers in the net?

• Which extensions of existing standards have to be made to represent the semantic meaning of formulae in historical documents?